# Humanoid Robots and Humanoid AI: Review, Perspectives and Directions

### v2

@@ 4.5.1. Humanoid Digital Twins and Metaverse
First, humanoid digital twins and metaverse surpass humanoid capabilities and capacity. Metaverse, digital twin, and virtual, augmented and mixed reality techniques can assist human-humanoid interaction, interfacing and collaboration in creating metaverse-humanoid-human collaborative design, development, workplace and tasking settings (Suzuki et al., 2022; Chang and Hayes, 2020; Baratta et al., 2024; Walker et al., 2023). On-humanoid, on-body or on-environment animation, metaverse and mixed reality can augment humanoid’s collaborative programming, teleoperation in projected environment and background, navigation trajectory visualization, immersive and expressive perception and telepresence. For example, by creating a 3D presentation of a humanoid and its workplace, a humanoid can be operated, monitored or tutored for specific tasks in its working environment (Ishiguro et al., 2024). Instructors with multi-modal virtual, augmented and mixed reality devices can operate and control humanoids and conduct various tasks, such as imposing humanoid movement restrictions, trajectory teaching, changing humanoid joint angles, and creating task-specific programs. For human-unfriendly environments, humanoids can be teleoperated to undertake specific tasks remotely using projection mapping to facilitate robotic tasks (Darvish et al., 2023).
@@ 4.5.2. Humanoid Animation, Imitation and Simulation to Operation
Humanoid Imitation to Operation Imitation learning is another type of simulation or demonstration. Humanoids learn by demonstrations or are taught by a supervisor, a demonstrator or examples, such as from virtual simulations or a filmed physical scenario (a video), and then clone, imitate or adjust the demonstrations or examples. Humanoids may also learn an underlying optimal reward function or policies from teachers, adapt to taught examples, develop or generate new behaviors (Tagliabue and How, 2024). Humanoid imitation thus supports behavior cloning, intent learning, behavior or scene prediction, recommendation or generation. By integrating techniques like DeepFake, animation, LLMs and digital twins, humanoid imitation could learn from sequential, hierarchical, multimodal and live human activities, and social activities; processes, mechanisms and strategies for handling complex tasks; general and specific skills for problem-solving; and intention, morality and humanity exhibited during human actions and interaction. Various extensions include learning by imitation, learning to imitate (Sun et al., 2023), generation by imitation, domain adaptation or domain transfer. Humanoid imitation also face issues like imitation-to-reality gaps, in particular, for unseen, new and open world objects, scenarios and tasks, and efficient and effective in-distribution, in-domain and in-context imitation and imitation at scale.
@@ 4.5.3. Humanoid Demonstration to Generation
Humanoid Demonstration-to-Operation Humanoid demonstration illustrates robots with expected functions, tasks, scenarios, behaviors, actions, and processes. Humanoids can then learn from single, few or sequential demonstrations to operate, plan, navigate or evolve over existing, new, unseen, open or lifelong tasks or environments (Argall et al., 2009). Humanoids may learn from human demonstrations (Brohan et al., 2023b), such as positioning, posture, behaving, emotional, vocal and facial expressions, and ethical responses, to improve, evolve or develop their world knowledge, human-like actions and responses. Humanoids may also clone, optimize and expand the learned knowledge, activities or intelligence from demonstrations and further forecast and generate new actions, knowledge or intelligence through generative AI. Humanoid learning by demonstration leverages humanoid simulation and mitigates sim-to-real gaps by enhancing humanoid’s familiarity and realisticness of the world and building robot experience, memory and reasoning. Similar to humanoid simulation, humanoid demonstration faces demonstration-to-reality gaps (Demo2Real gaps for short). Such gaps may be caused by demonstration biases, distribution shift, misalignment, or the limitation and insufficiency of demonstrations (e.g., from zero shot, one shot to few shots), settings, capabilities, and capacity. Humanoid learning by demonstration still faces significant challenges for unstructured, ill-structured, unseen-structured, unknown-structured, transitional, lifelong and nonstationary settings and environments. Both humanoid simulation and demonstration are also challenged by zero, new, evolving, unseen, unknown and open objects, contexts, scenarios or tasks.
@@
Humanoid Generation-to-Operation With the fast-pace development of GenAI including LLMs, LMMs, VQA, VLPs, VLMs and VLAs, automated generation becomes feasible and essential in simulating and producing AI and realizing human-level intelligence, substantially improving the intelligence of humanoids. Humanoid intelligence generation may present new opportunities and challenges than existing LLMs for image-to-text, video-to-text, text-to-image, and text-to-video, making the generation suitable for humanoid settings. Examples are 1) single-modal generation, e.g., generating humanoid behaviors, actions, responses, and emotions; and 2) cross-modal generation, such as image-to-speech, image-to-conversation, image-to-emotion, animation-to-speech, animation-to-conversation, animation-to-singing, animation-to-dancing, text-to-speech, text-to-conversation, text-to-action, and text-to-emotion (Brohan et al., 2023b; Meng et al., 2024). Humanoid generation could be built on humanoid animation, simulation, imitation and demonstration, where animation, simulation, imitation and demonstration serve as supervision, providing one to few shots for robot learning and further generation. Zero-to-few shot learning, weakly-supervised learning, transfer learning, contrastive learning, and Knowledge distillation would be useful to augment humanoid learning and generation by LLMs or LMMs.
@@ Acknowledgements.
Appendix A List of Terminology
Below, Table 1 lists the key terminology used in the literature and in this paper. Here, their statements are more general, supplementing those discussed in the paper.
Table 1. List of terminology.
Term Statement Relevant terms Examples
AI humanoids Humanoid robots driven by AI and holding certain human intelligence Humanoid AI robots, AI humanoid robots, AI-driven humanoid robots, intelligent humanoids Ameca, Atlas, Geminoid, Pepper and Sophia
AI robots A robot with certain AI facilities, features or functions, empowered by AI technologies AI-powered robots, AI-enabled robots H1, GR-1, Figure 01, Optimus
AI robotics A research area combining robotics and AI to make robots intelligent AI robots, humanoid AI robotics, humanoid AI robotics
Human intelligence Intellectual capabilities of humans, enabling human functions and behaviors, etc. Human intelligence, mental quality
Human qualities Characteristics and capabilities distinguishing humans from other beings, such as cognitive, emotional, and social aspects Human traits, human attributes, human characteristics
Humane humanoids Humanoids demonstrating human qualities (such as being compassionate, kind, and sympathetic) in treating living beings Human-like humanoids, human-level humanoids
Humanoid AI A symbiotic subfield of AI to enable intelligent humanoid robotics, synthesizing disciplines and areas including robotics, AI, human science, and social science Humanoid AI research, Humanoid AI studies
Humanoid appearance The observable physical characteristics of humanoids Human appearance, human phenotype, human form, human look Optimus, Sophia, Ameca
Humanoid behaviors Actions and reactions undertaken by humanoids, such as hand and leg movement, posture, and gesture Human behaviors, human movement, human poster, human gesture Atlas, H1
Humanoid functions Biological, psychological and social capacities and activities of humanoids Human functions, human roles, human features
Humanoid humanity The qualities (being humane, compassionate, and sympathetic, etc.) that make humanoids similar as humans Human-like humanoid qualities
Humanoid intelligence The intellectual characteristics and capabilities of humanoids, typically replicating human intelligence
Humanoid learning Learning mechanisms, capabilities and activities in humanoids Humanoid reasoning, humanoid inference
Humanoid robots Robots replicating human appearances, functions, behaviors or qualities in full or in part for humanoids Androids, mechanical men, and anthropomorphic robots
Humanoid senses Replicating physiological and mental human senses with world perception and sensory modeling on humanoids Visual, auditory, gustatory, olfactory, and cognitive senses
Humanoid structures Replicating human structures with human-looking heads or bodies of humanoids
Human-level The ability to perform tasks at the level of human competence Human-level AI, human-level robots, human-level humanoid robots
Human-like Analogizing (anthropomorphism, not necessarily identical) human characteristics and qualities, such as appearances, structures, functions, behaviors, or other human qualities Human-like AI, human-like robots, human-like humanoid robots Ameca, Geminoid
Human-looking The resemblance (not necessarily identical) of observable human appearances, including general forms, structures, features or behaviors Human-looking robots, human-looking humanoid robots Pepper, Sophia
Uncanny valley effect The phenomenon of becoming almost but not quite human, transferring positive to negative feelings Uncanny valley phenomenon, Uncanny valley hypothesis
Appendix B Perspectives of Actionable Humanoids
Addressing the gaps of existing humanoids by technical advances in Section 4, we spotlight perspectives of real-world humanoids: empowering humanoid actionability with AI advances, enabling live humanoids in complex contexts, decentralizing humanoids with on-robot, edge and cloud intelligence, and assuring humanoid quality of services.
B.1. Empowering Actionable Humanoid with AI Advances
The fast-paced evolution of GenAI to AGI will substantially transform and revolutionalize humanoid design paradigms. It has never been so promising to advance humanoids with human-like actionable intelligence (Cao, 2022a; Cao et al., 2007; Cao, 2018). The advances and futures of generative-to-human-level AI landscape can deepen and widen human-like humanoid studies towards new-form and new-path problems, research methods, and functional and practical means, fostering human-level humanoids. Human-level humanoids expect to handle various goals, functions, tasks, and applications. We highlight a few below: multi-task and multi-party settings; reinforcement from live feedback, refinement and adaptation; and omnimodal vision-language-action modeling and mind-to-action translation.
Multi-task and multi-party settings: Human-level humanoids expect to handle multiple tasks concurrently, recurrently or sequentially in an individual, team or adversarial mode (Brohan et al., 2023b). Examples of multi-task settings are hybridizing language parsing, vision tasks with control; dialogue-based tasking with coordination through exchanging information; and translating and reasoning tasks from perception or through conversations in natural language. Multi-party settings involve multiple humanoids, and humanoids with humans or with third parties, where they coordinate, collaborate or compete with each other to undertake tasks (Li et al., 2024). These multi-task and multi-party settings require corresponding design, coordination, communication, and control. Their use cases include identifying, recognizing and differentiating visual, vocal and textual information from different humans in a multi-party humanoid-human conversation, where the humanoid needs to respond correspondingly to the right party in their interest and preferred sentiment.
Reinforcement from live feedback, refinement and adaptation: In live scenarios, humanoids must acquire live feedback from humans and adapt to online interaction, coordination, communication, tasking and teaming scenes. Real-time adaptation to live workflows, dynamics and changes forms essential functions yet challenges to limited on-humanoid intelligence. In such live settings, integrating on-humanoid, edge and cloud intelligent services, learning, reinforcement and fine-tuning would be essential (Cao, 2022b; Wang and Zhang, 2024). For example, a humanoid with feedback-based tasking and coordination can receive, process and transform human feedback on task performance, validate its responses, and iteratively fine-tune actions and responses over feedback. To enable live feedback-based humanoid refinement and adaptation, a humanoid should incorporate RLHF into its online modeling by representing feedback and interaction and generating live responses. It then iteratively optimizes objectives, refreshes functions, and updates responses over tasking.
Omnimodal vision-language-action modeling and mind-to-action translation: as discussed in Sections 4.3 and 4.4, transitional omnimodal modeling adapts humanoids to the settings and function requirements of real-time, interactive multimodal fusion, translation and transformation of multiple modalities of input signals. For example, a humanoid may upgrade LLMs and VLP from language-image pretraining and language-video pretraining to mind-to-action and perception-to-behavior modeling, in order to understand and reason about multimodal information fusion, and translate input recognition to decision actions. Transitional omnimodal humanoids can be empowered by Transformer (Vaswani et al., 2017), vision transformer (Li et al., 2022), generative pretrained LLMs, VLMs (Stone et al., 2023), VLA models (Brohan et al., 2023b) on the Internet-scale data. Humanoids can be empowered by LLMs and LMMs for real-time conversations, benefiting from their prompting mechanisms such as reinforcement learning with human feedback (RLHF) and chain-of-thought prompting to transform multi-step into intermediate reasoning (Wei et al., 2022). Humanoids can simultaneously fuse multiple sensory chancels for complex tasking, such as applying vision-language translation for real-time multi-lingual conversations, and undertaking multi-type tasking, multimodal interactions, and pathway-oriented processing.
B.2. Humanoid Live Tasking in Complex Environments
The ultimate goal of designing humanoids is to make them actionable in live settings for live tasking. Humanoid tasking in complex live scenes requires humanoid to operate in real-world settings and environments (Cao, 2015). Live environments may involve multiple robot sensors, multiple parties like humans and other robots, various data characteristics and complexities, low-quality backgrounds, etc. Live tasking may involve multiple goals and tasks and expect real-time processing and actions, and immediate response and feedback. This requires humanoids to be capable of diverse capabilities and intelligences, such as sensor and signal coordination and fusion, instant multimodal response generation and alignment, multi-party coordination, and human-humanoid-shared mental modeling.
Data characteristics and complexities-resilient humanoid tasking enables humanoids to work in poor quality environments and be agile and resilient to complex data and behaviors during tasking (Cao, 2018). Data may comprise historical and background information, such as from the Internet, live data and behaviors during interactions and processes, and external information required during operations. Making humanoid functions and operations responsive to relevant sources of data and their data changes in real time poses significant challenges and opportunities for efficient live humanoid learning and computing from complex data.
Real-time multi-party humanoid tasking facilitates tasks with multiple robots, humans, other agents or systems (e.g., metaverse, attackers or defenders). Moderating real-time humanoid-to-humanoid, human-humanoid, and human-humanoid-agent teaming, communication, interaction, collaboration, or competition becomes increasingly demanding for real-world robotic applications (Kanda et al., 2004; Li et al., 2024). It discloses many challenges and innovation opportunities, such as live multi-party humanoid-human dialogue, real-time multimodal and multi-sensor information fusion and alignment, and seamless cooperation between humanoids and human collaborators.
Bilateral human-humanoid-shared modeling. Existing humanoid studies often focus on one-sided manipulation, i.e., empowering humanoids towards intelligent capabilities and tasking. To make humanoid humane and human-like, bilateral human-humanoid-shared modeling would be essential, where humanoids are treated as pseudo human beings for more natural, human-like interactions with humans, while humans can also engage humanoids in natural ways. This requires new theories and tools, such as for bilateral understanding, interaction, conversation, identification, recognition, collaboration and alignment both from human to humanoid and from humanoid to human. This means both humans and humanoids should 1) have the abilities to identify, select and engage interested partners; 2) actively and proactively initiate, control and manage conversations, tasks and engagement with the counterparts on-demand; and 3) satisfy various conditions, rights and criteria in determining their interactions and collaborations.
B.3. Humanoid Open Tasking in Open Worlds
Live actionable humanoids confront open tasking requirements and open worlds. In fact, real-time interactive humanoids and human-humanoid systems in live environments are also open complex intelligent systems (Cao, 2015), living in open environments, undertaking open tasks, and acting under open settings.
Open humanoid systems. A live humanoid is open, immersing itself in various types and aspects of partial to systemic opennesses (Cao, 2015). Examples include internal interactions within and between objects and subsystems of a humanoid; external interactions between a humanoid and its community; dynamics of humanoid parts, subsystems and systems; and uncertainties of states, state transition, and performance of humanoid parts and system. This may trigger humanoid open structuring and automated architecture search in humanoid system analysis and design. Automated humanoids in complex environments require evolving, on-demand and self-organizing architecture search, structure configuration, modularization, model logic search, parameter discovery, memory customization, and modular and functional activation and deactivation. These will cultivate new designs and mechanisms to characterize, define, enable, organize, manage and evaluate open humanoid systems.
Open humanoid environments. The openness in an open environment for a humanoid may be embodied by various scenarios: 1) uncertain, changing, unseen and unknown contexts; 2) dynamic and uncertain information, energy and communication flow between humanoids and their environments; 3) changing constituents and structures of a humanoid; 4) evolving tasking requirements and capabilities of a humanoid; and 5) potential nonstationary quality, service and performance of a humanoid. Handling such open scenarios requires inbuilt dynamic, online, active, evolving and self-organizing humanoid capabilities, i.e., open capabilities, for (re)alignment, (re)cognition, (re)arrangement, (re)manipulation, adjustment, (re)adaptation, and transfer.
Open humanoid settings may broadly refer to 1) open input scenarios, such as open data, open domain, open vocabulary, open context and open environment, which are dynamic and evolving with change, drift and shift over time, space, frequency, structure/distribution or semantics; 2) open tasking requirements, as discussed below on open humanoid tasking; and 3) open output requirements, such as open set (with new and unseen labels), open-structured outputs (e.g., changing, new and unknown forms, formats, layouts, distributions), open evaluation (e.g., with evolving criteria and objectives), and open feedback (e.g., changing reviews, opinions, result preference, and quality expectation on outputs). These require humanoids to be sufficiently flexible, adaptive, evolving and self-organizing.
Open humanoid tasking. Live humanoid tasking involves open humanoid systems, environments and settings. It also confronts open tasking, e.g., open data processing (e.g., processing data with evolving, uncertain, unknown and unseen nature); open goal settings (e.g., without firm and predefined goals and objectives); open perception (such as perceiving open worlds and enabling evolving, reconfigurable and self-developing perception tools and capabilities); open representation (e.g., with embedding capabilities of coping with open inputs); open (world) learning (e.g., under open settings, with evolving learning objectives, changing learning structures, or open output requirements) (Kejriwal et al., 2024); open reasoning (e.g., reasoning over uncertain actions, without constraints, or under open conditions); open actioning (e.g., enabling new, changing and beyond-design actions and behaviors); open control (e.g., enabling control in open environment and evolving tasking); open planning (e.g., pursuing plans in open contexts, planning by learning from open data and objectives); open evaluation (e.g., changing or adaptive evaluation criteria, methods or measures over input conditions and human feedback change); open optimization (e.g., without predefined optimization objectives or criteria, evolving objectives over input condition change); and open domain adaptation (e.g., the target domains are evolving, uncertain, new or unseen). These nurture a new area of open tasking for humanoid robotics, AI.
B.4. Decentralized Humanoids with On-Humanoid, Edge and Cloud Intelligence
Like smart devices such as smart phones and electric vehicles, humanoids should host both self-sustained on-device intelligence and networked intelligence augmented through networking and communication with other intelligent facilities and agents. This implies decentralized humanoids integrating on-humanoid, edge and cloud intelligence.
Decentralized Humanoids Decentralized humanoids build on decentralized AI, integrating decentralized autonomous organizations, blockchain technologies, Web3, decentralized finance, cryptocurrencies, decentralized science, and other decentralized technologies into enabling AI tasks, behaviors and systems over the web, cloud, edge or on device (Cao, 2022b). Decentralized humanoids empower on-humanoid AI for on-robot live perception, action and response and peer-to-peer on-robot AI tasks and activities. They access edge networks with a humanoid community for data-, model- and capability-sharing and edge intelligence (Cao, 2024). Such edges link cloud facilities and learning systems, enabling humanoid cloud and augmenting humanoid intelligence, e.g. by cloud AI like ChatGPT. Humanoid edge intelligence enables humanoid interfacing, interaction, conversation, communication, cooperation, collaboration, sharing, storage, memorization, action and intervention leveraging on-humanoid intelligence. Humanoid edge networks support privacy-preserving, secure, trustful, responsible, and efficient operations, control, communication, and collaborations. Humanoids decompose, split, dispatch, distribute, pipeline, orchestrate, aggregate or balance their roles, responsibilities, tasks, data, and resources among the humanoid edge network.
On-Humanoid, Edge and Cloud Humanoids On-humanoid, edge and cloud Humanoids foster a new era of humanoid robotics with enormous new challenges and opportunities for robotics, humanoid and decentralized AI. Humanoids enabled by LLMs such as Aemca by ChatGPT are cloud humanoids (Engineered Arts, 2023). Cloud humanoids leverage their limited on-robot capabilities and capacity by live high-speed connection and communication to humanoid cloud platforms and enabling services, such as ChatGPT for performing LLMs. This makes cloud humanoids dependent on specific cloud infrastructure, computing and services, making it costly and inflexible for operations and development. It also makes humanoids restrictive to Wifi or telco network connection, making humanoids incapable of on-device automation and end-to-end solutions, causing issues including latency, unreliability and risk in real-time interactions, coordination and communications. Cloud humanoids may also raise concerns on sharing privacy, safety and security of humanoids and their tasks and data to cloud service providers, making regulation difficult to be accountable instantly.
Accordingly, on-humanoid and edge humanoid development emerge for more independent, automated decentralized robotic and humanoid intelligence. On-humanoid intelligence, applications and services require energy-efficient, information-compressed, real-time, personalized, interactive and privacy-preserving hardware, platforms, applications, and communications. A humanoid is an independent automated, self-organizing intelligent agent. It senses, interacts, navigates, judges, plans, acts, controls and decides things by itself, holds its own goals, undertakes actions, and assesses live environment independently in real time. These require on-humanoid perception, recognition, conversation, interaction, regulation, etc. For example, LLMs and VLAs trained on cloud will have to be converted to humanoid-compatible LLMs and VLAs, such as by selecting mobile apps-discriminate parameters through selective parameter-efficient fine tuning. Edge humanoids nurture edge intelligence across an edge humanoid network. Humanoid edge networks handle energy-efficient, privacy-preserving, humanoid-neutral and -specific data acquisition, recognition, analytics, learning, reasoning, planning, and communication. Humanoid edge networks facilitate the networking, communications and shared functions across the network and client humanoids.
Figure 4 illustrates a framework of decentralized humanoid AI systems, where humans interact with humanoids through multimodal channels and manipulate humanoids through humanoid edge desktops. Humanoids connect to their cloud servers and communicate the perceived information with humans to the server (Engineered Arts, 2023; Li et al., 2024). The cloud servers compute and learn the information, retrieve, request or execute Internet-based third-party cloud services such as ChatGPT and GitHub-based open source APIs or applications. The processed responses from the cloud servers are communicated to client humanoids for executing corresponding actions.
Figure 4. Decentralized humanoids: On-humanoid, edge and cloud integrative humanoid AI framework.
B.5. Assuring Humanoid Quality of Services
In Section 3.2, we have discussed various issues relating to the quality of services (QoS) of humanoids from functional and nonfunctional perspectives. Here, we further expand the discussion on QoS to humanoid design quality, implementation quality, service quality, and output trust. These are important for making humanoids mindful, actionable and ethical.
Addressing biases, fairness and trust in humanoids. Humanoids may implicate various biases (Ferrara, 2023) such as sampling, modeling, and evaluation biases. Sampling biases may further generate biases and fairness issues, such as sample set biases, distributional biases, where samples may be associated with demographic, confirmation, memory, linguistic, ethnic/cultural, political or ideological biases. Modeling biases may be caused by hypothesis biases, methodological biases, architectural biases, input-model or model-task matching and fitting biases. Evaluation biases appear in biased or unfair evaluation methodologies, objectives, objective functions, evaluation measures, and evaluation processes. Addressing such biases and unfairnesses demand humanoid trust, fostering trustworthiness of humanoid designs, operations and actions from various aspects, such as faith, fear, confidence, feeling, valence, control and reciprocation of humanoids. It may also require establishing the corresponding humanoid norms, regulations, and laws (Cho et al., 2015).
Assuring humanoid design quality. GenAI and LLMs models exhibit QoS issues such as hallucination, and biased, superficial, irrelevant, too general and non-personalized responses (Huang et al., 2023). Their outputs may also present low actionability, ethical, security and privacy concerns (Cao, 2022a). These issues are grounded on insufficient design, also troubling decentralized humanoids enabled by such LLMs and GenAI, making their outputs vulnerable, untrustworthy and nonactionable by users. Humanoid capability maturity must examine such issues by either improving their dependent LLMs and variants or addressing LLM issues within humanoid designs. Here are a few directions of assuring humanoid design quality: conducting energy-efficient and parameter-efficient finetuning of pretrained LLMs or their results; incorporating LLM results as prior knowledge and training vertical robotic models; developing algorithmic debiasing and mitigating biases in training sampling; and performing active and online refinement of pretrained models, and human-in-the-loop refinement and decision-making. This demands bias and trust resilient design, modeling, learning, optimization, evaluation and feedback theories and techniques to manage complex biased scenarios, design generalization, bias and error propagation, unanticipated capability or result emergence, and alignment with human feedback for humanoids.
Assuring humanoid implementation quality. Humanoid implementation quality is attributed to decomposed quality of input and data, processing, hardware, software, interfacing, and output respectively (Ishiguro et al., 2024). First, collecting historical, present and future inputs and real-time data streaming and recordings during humanoid operations and human-humanoid interactions, and their manipulation and processing require data and processing quality assurance. Data and processing quality assurance should address data quality issues such as biases, inconsistencies, irregularity, and noises, and ensure fair and unbiased data sampling, cleaning, transformation, augmentation, denoising, and debiasing to make data ready. Second, the humanoid quality assurance of hardware and software is specially fabricated to fulfill functions and performance measures in Section 3. Accordingly, their implementation, quality and performance require assurance such as sufficient matching between hardware and software, systematic integrity, function cohesion, coding biases, and privacy and security assurance. Last but not least, the humanoid quality assurance of interfacing and outcomes requires addressing issues such as fair and unbiased usability of user interaction, interfacing and support; interpretability and explainability of workflow, factors and outputs; auditability and accountability of operational processes, actions and results; monitoring, evaluating and intervening potential consequences, risk, safety, security or trust concerns of outputs; and overseeing, mitigating and regulating business, societal or human effects and impact.
Appendix C Comparison of Functional and Nonfunctional Specifications
Here, we further compare the functional and nonfunctional specifications following the discussion in Section 3.
Following the discussion in Section 3.1, Table LABEL:tab:functionalspecifications summarizes the areas and aspects of functional humanoids, and their general robotic functions, humanoid specialties, and enabling approaches.
Then, taking forward the discussion on nonfunctional specifications in Section 3.2, Table LABEL:tab:nonfunctionalspecifications summarizes their areas and aspects of nonfunctional specifications of humanoids, and their research questions, attributes and enabling approaches.
Appendix D Summary and Comparison of Humanoid Robots
Here, we further discuss 30 types of humanoid or relevant robots, which can be aligned with the review and perspective discussion in the paper.
Table LABEL:tab:humanoids summarizes and compares 30 humanoid robots reported in the literature or on market. Interested readers can find the most recent humanoids such as Ameca, Altas, Figure 01, GR-1, H1 and Optimus.

### v3

@@ 4.5.1. Humanoid Digital Twins and Metaverse
First, humanoid digital twins and metaverse surpass humanoid capabilities and capacity. Metaverse, digital twin, and virtual, augmented and mixed reality techniques can assist human-humanoid interaction, interfacing and collaboration in creating metaverse-humanoid-human collaborative design, development, workplace and tasking settings (Suzuki et al., 2022; Chang and Hayes, 2024; Baratta et al., 2024; Walker et al., 2023). On-humanoid, on-body or on-environment animation, metaverse and mixed reality can augment humanoid’s collaborative programming, teleoperation in projected environment and background, navigation trajectory visualization, immersive and expressive perception and telepresence. For example, by creating a 3D presentation of a humanoid and its workplace, a humanoid can be operated, monitored or tutored for specific tasks in its working environment (Ishiguro et al., 2024). Instructors with multi-modal virtual, augmented and mixed reality devices can operate and control humanoids and conduct various tasks, such as imposing humanoid movement restrictions, trajectory teaching, changing humanoid joint angles, and creating task-specific programs. For human-unfriendly environments, humanoids can be teleoperated to undertake specific tasks remotely using projection mapping to facilitate robotic tasks (Darvish et al., 2023).
@@ 4.5.2. Humanoid Animation, Imitation and Simulation to Operation
Humanoid Imitation to Operation Imitation learning is another type of simulation or demonstration. Humanoids learn by demonstrations or are taught by a supervisor, a demonstrator or examples, such as from virtual simulations or a filmed physical scenario (a video), and then clone, imitate or adjust the demonstrations or examples. Humanoids may also learn an underlying optimal reward function or policies from teachers, adapt to taught examples, develop or generate new behaviors (Tagliabue and How, 2024). Humanoid imitation thus supports behavior cloning, intent learning, behavior or scene prediction, recommendation or generation. By integrating techniques like DeepFake, animation, LLMs and digital twins, humanoid imitation could learn from sequential, hierarchical, multimodal and live human activities, and social activities; processes, mechanisms and strategies for handling complex tasks; general and specific skills for problem-solving; and intention, morality and humanity exhibited during human actions and interaction. Various extensions include learning by imitation, learning to imitate (Sun et al., 2024), generation by imitation, domain adaptation or domain transfer. Humanoid imitation also face issues like imitation-to-reality gaps, in particular, for unseen, new and open world objects, scenarios and tasks, and efficient and effective in-distribution, in-domain and in-context imitation and imitation at scale.
@@ 4.5.3. Humanoid Demonstration to Generation
Humanoid Demonstration-to-Operation Humanoid demonstration illustrates robots with expected functions, tasks, scenarios, behaviors, actions, and processes. Humanoids can then learn from single, few or sequential demonstrations to operate, plan, navigate or evolve over existing, new, unseen, open or lifelong tasks or environments (Argall et al., 2009). Humanoids may learn from human demonstrations (Zitkovich et al., 2023), such as positioning, posture, behaving, emotional, vocal and facial expressions, and ethical responses, to improve, evolve or develop their world knowledge, human-like actions and responses. Humanoids may also clone, optimize and expand the learned knowledge, activities or intelligence from demonstrations and further forecast and generate new actions, knowledge or intelligence through generative AI. Humanoid learning by demonstration leverages humanoid simulation and mitigates sim-to-real gaps by enhancing humanoid’s familiarity and realisticness of the world and building robot experience, memory and reasoning. Similar to humanoid simulation, humanoid demonstration faces demonstration-to-reality gaps (Demo2Real gaps for short). Such gaps may be caused by demonstration biases, distribution shift, misalignment, or the limitation and insufficiency of demonstrations (e.g., from zero shot, one shot to few shots), settings, capabilities, and capacity. Humanoid learning by demonstration still faces significant challenges for unstructured, ill-structured, unseen-structured, unknown-structured, transitional, lifelong and nonstationary settings and environments. Both humanoid simulation and demonstration are also challenged by zero, new, evolving, unseen, unknown and open objects, contexts, scenarios or tasks.
@@
Humanoid Generation-to-Operation With the fast-pace development of GenAI including LLMs, LMMs, VQA, VLPs, VLMs and VLAs, automated generation becomes feasible and essential in simulating and producing AI and realizing human-level intelligence, substantially improving the intelligence of humanoids. Humanoid intelligence generation may present new opportunities and challenges than existing LLMs for image-to-text, video-to-text, text-to-image, and text-to-video, making the generation suitable for humanoid settings. Examples are 1) single-modal generation, e.g., generating humanoid behaviors, actions, responses, and emotions; and 2) cross-modal generation, such as image-to-speech, image-to-conversation, image-to-emotion, animation-to-speech, animation-to-conversation, animation-to-singing, animation-to-dancing, text-to-speech, text-to-conversation, text-to-action, and text-to-emotion (Zitkovich et al., 2023; Meng et al., 2024). Humanoid generation could be built on humanoid animation, simulation, imitation and demonstration, where animation, simulation, imitation and demonstration serve as supervision, providing one to few shots for robot learning and further generation. Zero-to-few shot learning, weakly-supervised learning, transfer learning, contrastive learning, and Knowledge distillation would be useful to augment humanoid learning and generation by LLMs or LMMs.